OpenStudy (anonymous):

Rationalize 9/3 (sqrt 7) I have an idea, but I'm not sure if I am completely off course

OpenStudy (anonymous):

I assume you mean \[\frac{ 9 }{ 3\sqrt{7} }\]

OpenStudy (anonymous):

In that case you should multiply the top and bottom by \(\sqrt{7}\) so that the square root on the bottom disappears.